# Feigenbaum universality in subcritical Taylor-Couette flow

### v1

@@ 5 Prediction of the period doubling route to chaos
Following the seminal work of Feigenbaum (1979), who showcased the first comparison with experimental results, Feigenbaum (1980, 1982, 1983) further extended renormalisation theory to include the universal scaling of all periodic points, not limited to the central branch, at different values of the parameter. The so-called trajectory scaling function theory, which unravels the scaling between orbits featuring the same stability properties (i.e., having the same value of the multiplier), was later validated against experimental data on natural convection by Belmonte et al. (1988).
The occurrence of universality entails that, for any given unimodal map, one should be able to predict the entire structure of the bifurcation diagram near the accumulation point using data from only the initial few period-doubling bifurcations. Unfortunately, the method of trajectory scaling function does not provide an easy way to predict the bifurcation cascade. Moreover, orbits become aperiodic beyond the accumulation point, hindering direct comparison with orbits at different parameter values. In this section, we devise a method for predicting the unfolding of the bifurcation cascade up to and beyond the accumulation point.
@@
and provided the constants $\gamma$, $X$, $\mu$, and $a_{\infty}$ are suitably chosen. The derivation of (26) and (27) is given in Appendix A. To simplify the notation, we use functional powers to denote repeated application of a function (e.g. $f^{2}_{a}(x)=f_{a}(f_{a}(x))$). Also, we have dropped the F subindex when referring to Feigenbaum’s first ($\delta$) and second ($\alpha$) constants.
@@
If the four (map-dependent) parameters, $\gamma$, $X$, $\mu$, and $a_{\infty}$, can be estimated from the first few period doublings, then (26) provides a useful tool for predicting the behaviour of the dynamical system (25), subject to the constraints outlined by (27). In practice, fixing $n=2$ and using data up to PDn+2 (i.e. PD4) yields already reasonably accurate results. The procedure for prediction is straightforward, as we shall see shortly. All it takes is extracting template branches from the bifurcation diagram of the one-dimesnional dynamical system
@@ 6 Conclusions
We have studied in detail a period-doubling cascade that arises in subcritical counter-rotating Taylor-Couette flow employing small computational domains of annular-parallelogram shape. The cascade is seeded from a family of drifting rotating waves discovered by W22 and eventually leads to a chaotic regime fueled by the self-sustained process of wall-bounded shear flows.
A pivotal outcome of our analysis is the confirmation of Feigenbaum universality to an unprecedented degree of accuracy in the context of fluid dynamics. Our success is of course largely ascribable to the fast development of computational power over the past few decades, along with the outstanding numerical accuracy of the methods we employ to solve Navier-Stokes flows, which altogether has rendered thorough parameter sweeps feasible. That said, establishing universality has required very long and costly numerical simulations, a refined parametric exploration, and the deployment of very accurate numerical techniques for the detailed analysis of the asymptotic convergence of relative periodic orbits, the assessment of their stability and the estimation of the period-doubling bifurcation points. Determining both the first and second Feigenbaum constants to the third significant digit demands the precise computation of up to the seventh period-doubling bifurcation along the cascade. This has enabled an accurate estimation of the accumulation point and, with it, the unambiguous exposure of the self-similar structure of the period-doubling bifurcation cascade. A key element to make the analysis systematic has been the deployment of a convenient Poincaré section based on torque balance, which provides a robust sampling method to validate theories of unimodal discrete-time dynamical systems.
Furthermore, our results constitute the first confirmation of universality in a fluid flow problem subject to subcritical transition. It is also noteworthy that all solutions along the period doubling cascade exhibit spatial drift, a property that is unique to our system among the many previous studies on the subject (see table 1). As expected, the conventional Feigenbaum theory becomes applicable to the Taylor-Couette phase space once reduced by the method of slices (Budanur et al., 2015).
Drawing from Feigenbaum’s theory, we have further developed a method to predict the bifurcation diagram of the period-doubling cascade all the way up to the accumulation point and, remarkably, also beyond. This approach has broader applicability than the method of trajectory scaling functions (Feigenbaum, 1983; Belmonte et al., 1988), which is limited to scrutinising stable orbits with the same value of the multiplier. Our prediction method is simple: it only requires stretching and shifting a template bifurcation diagram produced by any smooth unimodal map at hand. Of course, the scaling and translating parameters must be chosen such that the template bifurcation diagram aligns with the period-doubling cascade of the target system (in our case, Taylor-Couette flow). This can be accomplished, for example, by matching two bifurcation points across systems. The remainder of the transformed template diagram provides the prediction for the target system. The prediction results are useful, among other things, to anticipate the location of periodic windows in the bifurcation diagram of the target system beyond the accumulation point. The emergence of a stable orbit in the Taylor-Couette system precisely within the expected parameter range attests to the exceptional predictive capability of the method.
@@
Although the Navier-Stokes equations should formally be considered an infinite-dimensional dynamical system, the dynamics are often confined to finite-dimensional manifolds in phase space, sometimes designated as inertial manifolds (Temam, 1989). This low dimensionality has recently attracted significant attention (Ding et al., 2016; Haller et al., 2023), as it affords valuable insight into the underlying dynamics of the flow, sometimes leading to simplified models whose numerical simulation is computationally more tractable. Our results can be interpreted as an extreme case of simplification that ultimately leads to a one-dimensional map representations.
@@
The observation of Feigenbaum universality implies that the dynamics around the accumulation point of our Taylor-Couette setup can indeed be approximated by a nearly-one-dimensional discrete map when analysed on an appropriate Poincaré section. Hence, our finding provides a handy playground for testing in fluid dynamics the application of iconic theorems from the early days of chaos theory, such as Li-Yorke or Sharkovsky’s (Sharkovskii, 1964; Li & Yorke, 1975), as well as cycle expansion theory (Cvitanović et al., 2016). It is our intention to explore the significance and implications of these theories to the field of fluid dynamics in the near future.
@@
BW’s and RA’s research have been both funded by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101034413. This research was supported by the Australian Research Council Discovery Project DP230102188 and the Spanish MICINN grant No. PID2020-114043GB-I00.
The authors report no conflict of interest.
@@ Appendix A Derivation of a universal approximation to a unimodal map and the conditions for validity
where we have made explicit the effect of the renormalisation operator upon repeated application. In order to get back to the original map, we can use in (44) the obvious proposition that repeated application of the original and shifted maps is related by $F_{a}^{n}(y)=f_{a}^{n}(y+X)-X$ for any $n\in\mathbb{N}$. Combined with the definition of the universal function (28), the approximation
@@
Let us now consider what conditions are necessary for the approximation (45) to be valid, besides the obvious condition that $2^{n}$ must be large. The reliability of (45) depends on the degree of trust one can place on the approximations (35) and (41). It is important to note that, in renormalisation theory, comparisons of functions are conducted on magnified scales. That is, while (35) may be a good approximation on the scale where $y/\alpha^{n}=(x-X)/\alpha^{n}$ is not too large, it may not hold true on the original scale where $y\sim O(1)$ (mathematically, this can be easily understood by considering the case where $\varPsi$ is topologically conjugate to $g$). In (45), it suffices to consider $U_{M}(\xi)$ with the argument in the range $\xi\in[-1,1]$, which encompasses the entire period doubling cascade of (29). Hence the restriction $|(x-X)/\Gamma_{n}|<1$, which makes (35) only locally valid. The approximation (41) is, again, local. Moreover, in order to use the approximation (40) at level $k=n$, $|\delta^{n}(a-a_{\infty})|$ must necessarily be small, so we require $|\delta^{n}M|\ll 1$.
The approximation (45) must of course also hold for $U_{M}(\xi)$ itself, as it is unimodal in the range $\xi\in[-1,1]$. We can therefore set $n=m$, $x=\tilde{x}$, $a=\tilde{M}$, $X=0$, $a_{\infty}=0$, $\mu=\tilde{\mu}$, $\gamma=\tilde{\gamma}$ and $f_{a}=U_{\tilde{M}}$ in (45)-(46) to obtain

### v2

@@ 5 Prediction of the period doubling route to chaos
Following the seminal work of Feigenbaum (1979), who showcased the first comparison with experimental results, Feigenbaum (1980, 1982, 1983) further extended renormalisation theory to include the universal scaling of all periodic points, not limited to the central branch, at different values of the parameter. The trajectory scaling function theory, which unravels the scaling between orbits featuring the same stability properties (i.e., having the same multiplier), was later applied to interpret experimental data on natural convection by Belmonte et al. (1988).
The occurrence of universality entails that, for any given unimodal map, one should be able to predict the entire structure of the bifurcation diagram near the accumulation point using data from only the initial few period-doubling bifurcations. Unfortunately, the trajectory scaling function method does not provide an easy way to predict the bifurcation cascade. Moreover, orbits become aperiodic beyond the accumulation point, hindering direct comparison with orbits at different parameter values. In this section, we devise a method for predicting the unfolding of the bifurcation cascade up to and beyond the accumulation point.
@@
and provided the constants $\gamma$, $X$, $\mu$, and $a_{\infty}$ are suitably chosen. The derivation of (26) and (27) is given in Appendix A. To simplify the notation, we use functional powers to denote repeated application of a function (e.g. $f^{2}_{a}(x)=f_{a}(f_{a}(x))$). Also, we have dropped the F subscript when referring to Feigenbaum’s first ($\delta$) and second ($\alpha$) constants.
@@
If the four (map-dependent) parameters, $\gamma$, $X$, $\mu$, and $a_{\infty}$, can be computed from the first few period doublings, then (26) provides a useful tool for predicting the behaviour of the dynamical system (25), subject to the constraints outlined by (27). In practice, fixing $n=2$ and using data up to PDn+2 (i.e. PD4) yields already reasonably accurate results. The procedure for prediction is straightforward, as we shall see shortly. All it takes is extracting template branches from the bifurcation diagram of the one-dimesnional dynamical system
@@ 6 Conclusions
We have studied in detail a period-doubling cascade that arises in subcritical counter-rotating Taylor-Couette flow employing small computational domains of annular-parallelogram shape. The cascade is seeded from a family of drifting rotating waves discovered by W22 and eventually leads to a chaotic regime fueled by the self-sustained process of wall-bounded shear flows. Although the Navier-Stokes equations should formally be considered an infinite-dimensional dynamical system, the dynamics are often confined to finite-dimensional manifolds in phase space, sometimes designated as inertial manifolds (Temam, 1989). This low dimensionality has recently attracted significant attention (Ding et al., 2016; Haller et al., 2023), as it affords valuable insight into the underlying dynamics of the flow, sometimes leading to simplified models whose numerical simulation is computationally more tractable. Our results can be interpreted as an extreme case of simplification that ultimately leads to a one-dimensional map representation.
A pivotal outcome of our analysis is the confirmation of Feigenbaum universality to an unprecedented degree of accuracy in the context of fluid dynamics. Our success is largely ascribable to the fast development of computational power over the past few decades, along with the outstanding numerical accuracy of the methods we employ to solve Navier-Stokes flows, which altogether has rendered thorough parameter sweeps feasible. That said, establishing universality has required very long and costly numerical simulations, a refined parametric exploration, and the deployment of very accurate numerical techniques for the detailed analysis of the asymptotic convergence of relative periodic orbits, the assessment of their stability and the computation of period-doubling bifurcation points. Determining both the first and second Feigenbaum constants to the third significant digit demands the precise computation of up to the seventh period-doubling bifurcation along the cascade. This has enabled an accurate estimation of the accumulation point and, with it, the unambiguous exposure of the self-similar structure of the period-doubling bifurcation cascade. A key element to make the analysis systematic has been the deployment of a convenient Poincaré section based on torque balance, which provides a robust sampling method to validate theories of unimodal discrete-time dynamical systems.
Furthermore, our results constitute the first confirmation of universality in a fluid flow problem subject to subcritical turbulent transition. A detailed study of period doubling cascades is contingent on the existence and identification of a stable solution. In subcritical shear flows stable solutions are rare and can only occasionally be found by working with minimal flow units and/or symmetry subspaces. We were fortunate enough to find one such solution in W22 that happens to be at the origin of a period-doubling cascade. It is also noteworthy that all solutions along the period doubling cascade exhibit spatial drift, a property that is unique to our system among the many previous studies on the subject (see table 1). As expected, the conventional Feigenbaum theory becomes applicable to the Taylor-Couette phase space once reduced by the method of slices (Budanur et al., 2015).
Drawing from Feigenbaum’s theory, we have further developed a method to predict the bifurcation diagram of the period-doubling cascade all the way up to the accumulation point and, remarkably, also beyond. This approach has broader applicability than the method of trajectory scaling functions (Feigenbaum, 1983; Belmonte et al., 1988), which is limited to scrutinising stable orbits with the same value of the multiplier. Our prediction method is simple: it only requires stretching and shifting a template bifurcation diagram produced by any smooth unimodal map at hand. The scaling and translating parameters must be chosen such that the template bifurcation diagram aligns with the period-doubling cascade of the target system (in our case, Taylor-Couette flow). This can be accomplished, for example, by matching two bifurcation points across systems. The remainder of the transformed template diagram provides the prediction for the target system. The prediction results are useful, among other things, to anticipate the location of periodic windows in the bifurcation diagram of the target system beyond the accumulation point. The emergence of a stable orbit in the Taylor-Couette system precisely within the expected parameter range attests to the exceptional predictive capability of the method.
Our analysis further provides a theoretical explanation for the self-similarity that occurs in the reverse cascade following the accumulation point (see the appendix). Since the theoretical result holds universally for unimodal maps, the excellent agreement between our DNS data and the reference one-dimensional map implies that self-similarity can be reasonably expected to extend beyond the accumulation point also for the Navier-Stokes system.
@@
The observation of Feigenbaum universality in our Taylor-Couette setup provides evidence that the dynamics around the accumulation point can indeed be approximated by a nearly-one-dimensional discrete map when analysed on an appropriate Poincaré section. Our finding further provides a handy playground for testing the application of iconic theorems from the early days of chaos theory, such as Li-Yorke or Sharkovsky’s (Sharkovskii, 1964; Li & Yorke, 1975), as well as cycle expansion theory (Cvitanović et al., 2016), to a fluid dynamics problem. It is our intention to explore the significance and implications of these theories to the field of fluid dynamics in the near future.
As briefly commented in section 1, understanding the laminar-turbulent pattern-formation aspects of subcritical transition using stochastic theories (Hof, 2023) requires the prior occurrence of chaotic dynamics in the system. In shear flows, these incipient chaotic dynamics are only transient, such that the underlying chaotic saddle is not accessible through experiments or DNS. The scenario we have dissected offers a possible mechanism for the formation of one such chaotic saddle because our period doubling cascade (i) occurs globally in the full annulus, and (ii) is unstable to subharmonic (spatially modulational) perturbations, as can be inferred from W22. Both the full and elongated domains of figure 1b are simultaneously compatible with spiral turbulence and with the period doubling cascade we address here. Evidence for (ii) can indeed be drawn from W22, where the subharmonic instability of DRW, along with all solutions bifurcated from it in our small domain, was shown to contribute decisively to the formation of the laminar-turbulent helical pattern that is characteristic of the spiral turbulence regime. Therefore, although other mechanisms have been shown to produce localised chaotic sets that may be held responsible for intermittency in some cases (see Paranjape et al., 2023, for an example in channel flow), also a period doubling cascade might work concurrently with a modulational instability to the same effect.
[…]
@@
[Acknowledgements] This research is supported by the Australian Research Council Discovery Project DP230102188 and the Ministerio de Ciencia, Innovación y Universidades (Agencia Estatal de Investigación, project nos. PID2020-114043GB-I00 (MCIN/AEI/10.13039/501100011033) and PID2023-150029NB-I00 (MCIN/AEI/10.13039 /501100011033/FEDER, UE). BW’s and RA’s research has been funded by the European Union’s Horizon 2020 research and innovation programme (Marie Skłodowska-Curie Grant Agreement No. 101034413). RA has also been funded by the Austrian Science Fund (FWF) 10.55776/ESP1481224.
[…]
[Declaration of Interests] The authors report no conflict of interest.
[…]
[Author ORCIDs]B. Wang, https://orcid.org/$0$-0002-6229-0336; R. Ayats, https://orcid.org/$0$-0001-6572-0621; K. Deguchi, https://orcid.org/$0$-0002-3709-3242; A. Meseguer, https://orcid.org/$0$-0002-2022-2001; F. Mellibovsky, https://orcid.org/$0$-0003-0497-9052
@@ Appendix A Derivation of a universal approximation to a unimodal map and the conditions for validity
where we have made explicit the effect of the renormalisation operator upon repeated application. In order to get back to the original map, we can use in (44) the easily verifiable proposition that repeated application of the original and shifted maps is related by $F_{a}^{n}(y)=f_{a}^{n}(y+X)-X$ for any $n\in\mathbb{N}$. Combined with the definition of the universal function (28), the approximation
@@
Let us now consider what conditions are necessary for the approximation (45) to be valid, besides the one requiring that $2^{n}$ must be large. The reliability of (45) depends on the degree of trust one can place on the approximations (35) and (41). It is important to note that, in renormalisation theory, comparisons of functions are conducted on magnified scales. That is, while (35) may be a good approximation on the scale where $y/\alpha^{n}=(x-X)/\alpha^{n}$ is not too large, it may not hold true on the original scale where $y\sim O(1)$ (mathematically, this can be easily understood by considering the case where $\varPsi$ is topologically conjugate to $g$). In (45), it suffices to consider $U_{M}(\xi)$ with the argument in the range $\xi\in[-1,1]$, which encompasses the entire period doubling cascade of (29). Hence the restriction $|(x-X)/\Gamma_{n}|<1$, which makes (35) only locally valid. The approximation (41) is, again, local. Moreover, in order to use the approximation (40) at level $k=n$, $|\delta^{n}(a-a_{\infty})|$ must necessarily be small, so we require $|\delta^{n}M|\ll 1$.
The approximation (45) must also hold for $U_{M}(\xi)$ itself, as it is unimodal in the range $\xi\in[-1,1]$. We can therefore set $n=m$, $x=\tilde{x}$, $a=\tilde{M}$, $X=0$, $a_{\infty}=0$, $\mu=\tilde{\mu}$, $\gamma=\tilde{\gamma}$ and $f_{a}=U_{\tilde{M}}$ in (45)-(46) to obtain